# Seeking Certainty In Uncertainty: Dual-Stage Unified Framework Solving Uncertainty in Dynamic Facial Expression Recognition

### v1

@@ 2.1. Dynamic Facial Expression Recognition
In the realm of deep learning, foundational work utilizing 3D Convolutional Neural Networks (C3D) (Tran et al., 2015) has been instrumental in the DFER domain, enabling the simultaneous capture of spatial and temporal information (Tran et al., 2015). The subsequent integration of RNNs (Li et al., 2019) and LSTMs (Singh and Nasoz, 2020) has markedly enhanced the ability to model temporal dynamics. These advancements have been complemented by traditional feature extraction techniques employing architectures such as ResNet (He et al., 2016) and VGG (Simonyan and Zisserman, 2014). Methods such as Freq-HD (Tao et al., 2023) and NR-DFERNet (Li et al., 2022) focus on selecting frames with emotions. More recently, the adaptation of Transformer architectures from the NLP sector to the visual domain, through ViT (Dosovitskiy et al., 2020), has introduced a novel approach for feature extraction from images (Sun et al., 2023; Chen et al., 2023b; Sun et al., 2024). And methods (Tao et al., 2024; Li et al., 2023b; Zhao and Patras, 2023; Foteinopoulou and Patras, 2023) based on CLIP-ViT have been proposed to address the DFER task due to CLIP’s powerful prior knowledge. The self-attention mechanism inherent in ViTs offers improved efficacy in discerning relevant information from images for DFER tasks. Moreover, the introduction of Former-DFER (Zhao and Liu, 2021), which utilizes both spatial and temporal transformers, has demonstrated commendable performance, further substantiating the potential of transformer-based models in addressing DFER challenges.
@@ 2.2. Learning with Uncertainty
Learning with Noisy Label. Uncertainty refers to annotation ambiguity in FER task, which belongs to noisy label. The issue of learning with noisy label has garnered significant interest in the machine learning community in recent years (Zhang et al., 2022; Tan et al., 2021; Feng et al., 2023b; Tu et al., 2023). This surge in attention is largely due to the recognition that noise is an inevitable aspect of the dataset labeling process (Zhang et al., 2022; Li and Sun, 2022). Initial approaches were primarily focused on Sample Selection, a method for pruning data that contains noise (Feng et al., 2023a; Bai et al., 2021; Cheng et al., 2022). Techniques such as utilizing small-loss samples for training (Jiang et al., 2018; Han et al., 2018; Cheng et al., 2022; Yu et al., 2019)—for instance, MentorNet (Jiang et al., 2018) employs the small loss principle for the teacher model to select clean samples for the student model. Alternatively, some strategies involve adopting robust loss functions to mitigate the learning of noise, such as Mean Absolute Error (MAE) loss (Ghosh et al., 2017) and Symmetric Cross Entropy (SCE) loss (Wang et al., 2019). Methods like SELF (Nguyen et al., 2020) and SFT (Wei et al., 2022) rely on predictions from previous iterations to discern clean samples.

### v2

@@ 2.1. Dynamic Facial Expression Recognition
In the realm of deep learning, foundational work utilizing 3D Convolutional Neural Networks (C3D) (Tran et al., 2015) has been instrumental in the DFER domain, enabling the simultaneous capture of spatial and temporal information (Tran et al., 2015). The subsequent integration of RNNs (Li et al., 2019) and LSTMs (Singh and Nasoz, 2020) has markedly enhanced the ability to model temporal dynamics. These advancements have been complemented by traditional feature extraction techniques employing architectures such as ResNet (He et al., 2016) and VGG (Simonyan and Zisserman, 2014). Methods such as Freq-HD (Tao et al., 2023) and NR-DFERNet (Li et al., 2022) focus on selecting frames with emotions. More recently, the adaptation of Transformer architectures from the NLP sector to the visual domain, through ViT (Dosovitskiy et al., 2020), has introduced a novel approach for feature extraction from images (Sun et al., 2023; Chen et al., 2023b; Sun et al., 2024; li2023intensity; lee2023frame). And methods (Tao et al., 2024; Li et al., 2023b; Zhao and Patras, 2023; Foteinopoulou and Patras, 2023) based on CLIP-ViT have been proposed to address the DFER task due to CLIP’s powerful prior knowledge. The self-attention mechanism inherent in ViTs offers improved efficacy in discerning relevant information from images for DFER tasks. Moreover, the introduction of Former-DFER (Zhao and Liu, 2021), which utilizes both spatial and temporal transformers, has demonstrated commendable performance, further substantiating the potential of transformer-based models in addressing DFER challenges.
@@ 2.2. Learning with Uncertainty
Learning with Noisy Label. Uncertainty refers to annotation ambiguity in FER task, which belongs to noisy label. The issue of learning with noisy label has garnered significant interest in the machine learning community in recent years (Zhang et al., 2022; Tan et al., 2021; Feng et al., 2023b; Tu et al., 2023; ma2020normalized; liu2015classification). This surge in attention is largely due to the recognition that noise is an inevitable aspect of the dataset labeling process (Zhang et al., 2022; Li and Sun, 2022). Initial approaches were primarily focused on Sample Selection, a method for pruning data that contains noise (Feng et al., 2023a; Bai et al., 2021; Cheng et al., 2022). Techniques such as utilizing small-loss samples for training (Jiang et al., 2018; Han et al., 2018; Cheng et al., 2022; Yu et al., 2019)—for instance, MentorNet (Jiang et al., 2018) employs the small loss principle for the teacher model to select clean samples for the student model. Alternatively, some strategies involve adopting robust loss functions to mitigate the learning of noise, such as Mean Absolute Error (MAE) loss (Ghosh et al., 2017) and Symmetric Cross Entropy (SCE) loss (Wang et al., 2019). Methods like SELF (Nguyen et al., 2020) and SFT (Wei et al., 2022) rely on predictions from previous iterations to discern clean samples.
@@ 3.2. Coarse-Grained Pruning Stage
Explaining weights and pruning mechanisms from the perspective of gradients. The model assigns higher weights $w_{i}$ to accurately predicted samples because these samples typically have clear and reliable feature representations, resulting in lower loss. Specifically, the predicted values $f(x_{i},\theta)$ for these samples are closer to the true labels $y$, thereby reducing the error term in the loss function. The gradient of the WCE Loss is calculated as follows:
$\frac{\partial\mathcal{L}_{\text{WCE}}}{\partial(w_{i}f(x_{i};\theta))}=-\frac{1}{w_{i}f(x_{i};\theta)}(y-w_{i}f(x_{i};\theta))$
where $y$ represents the true labels. Samples with higher weights have larger gradients, making their influence on the model parameter updates more significant, thereby enhancing the learning of these samples’ features. During backpropagation, the model adjusts the weights based on the loss feedback, favoring samples with lower losses by assigning them higher weights, as their predictions are closer to the true labels, providing more reliable supervision signals. The model can autonomously identify and assign weights to data of varying quality. Additionally, we implemented a threshold pruning mechanism to exclude samples with weighted predictions $w_{i}f(x_{i};\theta)$ below the threshold $\lambda$. This process aims to further reduce the impact of low-quality samples on model training. In the pruned dataset, only samples with higher weighted predictions are retained, while low-quality samples are assigned lower weights and excluded, minimizing their influence on the model’s learning process. Conversely, high-quality samples, due to their higher weights, have a more substantial impact on gradient computation, positively contributing to the model’s final performance. And we did some visualization in the ExperimentLABEL:experiment part to show that low-quality samples have lower weights than the high-quality samples.